OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21 Choose one answer. A. (2, 3, 5) B. (-2, 3, 5) C. (2, -3, 5) D. (2, 3, -5)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Okay so there are three equations. 1. x + 3y -z = 2 2. x - 2y + 3z = 7 3. x + 2y - 5z = -21 First you would have to subtract equation 1 from equation 2.

OpenStudy (anonymous):

x - 2y + 3z - x - 3y + z = 7 - 2 I'm not going to give you the answer straight away, but I will help you solve this. :p Try and solve the equation I put. ^^

OpenStudy (anonymous):

Are you there? lol

OpenStudy (anonymous):

-5y+4z=5

OpenStudy (anonymous):

@babixz101

OpenStudy (anonymous):

Yes, now remember that equation because we're gonna need it later! Now we have to subtract equation 1 from equation 3. x + 2y - 5z - x - 3y + z = 21 - 2 Tell me once you've got the answer.

OpenStudy (anonymous):

-y-4z=19

OpenStudy (anonymous):

@babixz101

OpenStudy (anonymous):

No I realized I made a mistake. It should've been -21 - 2!! Sorry, but I'll still count your answer because I messed up. It should be -23 instead of 19. My mistake.

OpenStudy (anonymous):

Remember that equation that I said we would need later? Well we need it now. We have to add that equation (-5y+4z=5) and this equation(-y-4z=19) together. So the whole equation would be -5y+4z-y-4z=5-23 Tell me once you found the answer. Btw, sorry it is taking a while to respond. My computer is a bit slow today.

OpenStudy (anonymous):

should be -y-4z=-23 not -y-4z=19. Sorry again!